OpenStudy (anonymous):

x f(x) 2 8 If f(x) = 6x + 2, which of the following values would complete the table shown? 0 and 1 64 and 70 16 and 64 14 and 50

OpenStudy (anonymous):

f(x) = 6x + 2 f(2) = 6(2) + 2 f(2) = 12 + 2 f(2) = 14 f(8) = 6(8) + 2 f(8) = 48 + 2 f(8) = 50 answer is : 14 and 50

OpenStudy (anonymous):

thnk u